to get to school every morning you must turn onto wembley road at a set of traffic lights. the traffic lights for wembley road has a cycle of green for 30 seconds, oraange for 10 second, than red for 20 seconds.

1. what is the probability that the red traffic light will be red as you approach the intersection
2. over 3 school weeks how many days would expect the traffic lights to be red as you approach the intersection
3.designed an experiment where you simulate the operation of the traffic light to record whether the light is red as you approach the intersection. explain what devices you will need to represent the outcomes
4.complete the experiment 10 times and calculate the experiment probabily that the traffic lights will be red as you approach the intersection.

working out is needed thank you

6.use the probabilities calcuated in part 4 and 5 to estimate the number of days that the traffic light will be red over 3 school weeks.
1
Expert's answer
2012-08-17T08:16:43-0400
1. what is the probability that the red traffic light will be red as you approach the intersection
Answer: P=20/60=1/3
2. over 3 school weeks how many days would expect the traffic lights to be red as you approach the intersection
Answer: 3weeks*5days*Probability=15*(1/3)= 5 days